OpenStudy (anonymous):

Can you give me an equation for this please? Two bike riders, X and Y both start at 2 pm riding toward each other from 40 km apart. X rides at 30 km/h and Y rides at 20 km/h. If they meet after t hours, find when and where they meet. Thanks !

OpenStudy (anonymous):

since the time of meeting equals time elapsed so, so let t=total time elapsed......therefore, 30t + 20t = 40 (D=S*T) 50t = 40 t = 4/5 hrs = 48 min..so they met at 2:48pm Now distance can be calculated plugging in value of time , D=30 *(4/5) =24 km

OpenStudy (anonymous):

thank you!

OpenStudy (anonymous):

pleasure to help ) anytime @Sabrinatheunicorn